### v1

@@ Abstract
Foundation Models (FMs) are large-scale, pre-trained AI systems that have revolutionized natural language processing and computer vision, and are now advancing geospatial analysis and Earth Observation (EO). They promise improved generalization across tasks, scalability, and efficient adaptation with minimal labeled data. However, despite the rapid proliferation of geospatial FMs, their real-world utility and alignment with global sustainability goals remain underexplored. We introduce SustainFM, a comprehensive benchmarking framework grounded in the 17 Sustainable Development Goals with extremely diverse tasks ranging from asset wealth prediction to environmental hazard detection. This study provides a rigorous, interdisciplinary assessment of geospatial FMs and offers critical insights into their role in attaining sustainability goals. Our findings show: (1) While not universally superior, FMs often outperform traditional approaches across diverse tasks and datasets. (2) Evaluating FMs should go beyond accuracy to include transferability, generalization, and energy efficiency as key criteria for their responsible use. (3) FMs enable scalable, SDG-grounded solutions, offering broad utility for tackling complex sustainability challenges. Critically, we advocate for a paradigm shift from model-centric development to impact-driven deployment, and emphasize metrics such as energy efficiency, robustness to domain shifts, and ethical considerations.
keywords:
@@
Foundation Models, Earth Observation, Sustainable Development Goals, Geospatial AI, Benchmarking
Introduction
The Sustainable Development Goals (SDGs) are a collection of 17 global goals set by the United Nations General Assembly in 2015. They aim to achieve a better and more sustainable future by addressing global challenges related to poverty, inequality, climate change, environmental degradation, peace, and justice. The SDGs offer a shared framework for countries, organizations, and individuals to advance an economically, socially, and environmentally sustainable world by 2030. Earth observation (EO) techniques significantly contribute to sustainable development by providing critical data and insights that support the achievement of SDGs [1, 2, 3]. EO contributes to global challenges related to poverty, inequality, climate change, environmental degradation, peace, and justice. From monitoring climate change, biodiversity, and water resources to enhancing food security, urban planning, and disaster resilience [4, 5, 6, 7], EO provides critical information for tracking environmental and socio-economic trends. Through supporting evidence-based policy and providing an enormous amount of up-to-date data, EO accelerates progress toward a more sustainable, equitable, and resilient future.
Exploiting EO data at scale requires automated tools; foundation models (FMs) offer a powerful solution for handling the diverse, multimodal spatio-temporal data typical of geospatial applications. FMs are large-scale artificial intelligence (AI) models that are pre-trained on extensive, diverse, and often multimodal datasets using self-supervised learning, which involves learning from raw, unlabeled data via proxy tasks, or weakly supervised learning, which utilizes noisy or incomplete labels [8]. FMs contain billions of parameters (e.g., Google’s SegGPT with over 3 billion and Meta’s DINOv2 with around 1 billion parameters), and leverage vast computational resources to learn detailed patterns and representations from massive datasets, which enables them to generalize across a wide range of downstream tasks. With this definition, they serve as a universal backbone to support a wide range of applications across domains such as natural language processing (NLP), computer vision, and (geo)scientific computing. The term foundation emphasizes their role in providing a pre-trained, transferable representation that can be fine-tuned or adapted for domain-specific applications. Geospatial FMs usually contain hundreds of millions of parameters and are pre-trained on vast airborne and spaceborne data. They can then be adapted to specific downstream tasks, e.g., land cover classification, change detection, or disaster response mapping, without requiring separate models to be trained from scratch for each task [9]. Key characteristics of FMs include: (1) Scalability through efficient, distributed training; (2) Generalization via transferable representations across tasks; and (3) Few-/Zero-shot Learning, enabling adaptation with little or no labeled data.
@@ Introduction
Since 2023, the EO community has witnessed an explosion in the development of FMs, evolving from single-modal [10, 11, 12] to multi-modal [13, 14, 15, 16] architectures that integrate diverse data sources such as optical imagery, SAR, and auxiliary geospatial data. These advancements aim to enable more comprehensive and robust downstream applications. However, the rapid growth in the number of FMs over such a short period has led to a saturation of new architectures, resembling the clear trajectory we witnessed in the EO community during the deep learning boom around 2016, often marked by incremental improvements and limited consideration of their societal, environmental, and economical impact. Despite advancements in computational efficiency and scalability, a fundamental gap remains in grounding model development with global sustainability objectives. To see the actual impact of FMs on real-world challenges, such as climate monitoring, resource management, and humanitarian applications, there needs to be a stronger focus on the deployment phase (downstream tasks), which is widely regarded as the ultimate goal of Responsible AI in EO [17].
@@
To address this, we must shift the emphasis from mere model development to the deployment phase, where FMs are evaluated based on their impact on complex socio-environmental contexts. Performance evaluation should go beyond traditional benchmarks and include real-world scenarios, social good, and ethical considerations in the deployment stage. Furthermore, the interdisciplinary views of both AI and domain experts are required to ensure that FMs are not only optimized for accuracy but also for their contribution to long-term societal benefits. Without this shift, we risk an unsustainable model development cycle with no real-world impacts.
@@ 1. Do we need FMs, or are traditional, purpose-specific deep learning models sufficient for EO tasks?
Through systematic evaluation across key criteria, including transferability, data efficiency, energy costs, and operational impact, we aim to provide a grounded, rigorous assessment that challenges assumptions and clarifies the actual added value, if any, of FMs in this domain. Ultimately, our goal is to contribute to global sustainability objectives and promote social good, a key aspect that has been undermined in many current geospatial FMs. To evaluate FMs in EO rigorously, we introduce SustainFM, a benchmark aligned with all 17 SDGs that provides a new dataset, leverages pretrained FMs, and evaluates them for global sustainability, as conveyed in Fig. 1. This dataset spans a wide range of geographies (over 200 countries and regions across continents, as shown in global distribution in Fig. 2), sensor types, and spatial resolutions (from 0.5 m to 30 m) to reflect the diversity and complexity of global sustainability challenges. SustainFM combines 16 real-world tasks (dedicated to the first 16 SDGs), including asset wealth prediction, child and women’s health estimation, solar farm and settlement mapping, and detecting industrial smoke, marine pollution, and conflict-related damage. These tasks cover key learning tasks such as regression, change detection, segmentation, and classification, and are tackled using a variety of sensors such as Landsat, Sentinel-1/2, VIIRS, Gaofen-2, Harmonized Landsat and Sentinel-2 (HLS), Google Earth, and PlanetScope. By encompassing domains like climate resilience, food security, infrastructure monitoring, and humanitarian response, SustainFM enables a comprehensive assessment of FMs across transferability, data efficiency, and real-world utility. By bringing together experts with diverse backgrounds, from domain scientists and remote sensing specialists to AI ethicists, machine learning researchers, and statisticians, we contribute to SDG-17 (Partnerships for the Goals) and offer critical yet comprehensive insights into whether FMs can contribute to global sustainability, moving beyond academic benchmarks to real-world impact.
@@
Empirical Evidences
We compare various geospatial FMs and traditional deep models using the results obtained on the SustainFM Benchmark. SustainFM includes data from six continents and incorporates data from a wide range of satellite sensors with spatial resolutions from 0.5 to 30 meters. To ensure task diversity, we include a variety of AI tasks such as regression (Reg.), semantic segmentation (Seg.), change detection (CD), and classification (Cls.), all of which are grounded in real-world applications, and includes a total of 723,555 image patches with patch sizes ranging from 120×120 to 1024×1024 to reflect reality. We provide detailed information about the building blocks of SustainFM, its diversity, and investigated models in Section 5, Methodology.
Table 1 presents the performance of various FMs (i.e., CROMA [18], RemoteCLIP [19], SoftCon [20], DOFA [21], GFM-Swin [22], Prithvi [23], SpectralGPT [10], ScaleMAE [24], SSL4EO-S12 [25], and SatlasNet [26]) and traditional deep models (i.e., Vision Transformer (ViT) [27] and Resnet-50 [28]) across the SustainFM tasks. We report performance using root mean square error (RMSE) for regression and mean F1 score (mF1) for classification, to capture both the prediction of continuous values and the identification of categorical classes.
Table 1: Performance of Foundation Models and Traditional Methods across SustainFM tasks. Metrics are RMSE (↓) or mF1 (↑) as appropriate. The best performance per column is shown in bold, and the second-best is underlined.
Model #1 #2 #3 #4 #5 #6 #7 #8 #9 #10 #11 #12 #13 #14 #15 #16
CROMA 1.39 84.10 35.01 2.75 2.10 84.94 96.52 69.78 93.10 0.99 91.92 73.87 89.66 70.73 87.50 96.71
[…]
GFM-Swin 1.45 87.74 34.74 2.71 2.08 87.89 94.86 75.83 93.59 1.04 93.78 78.05 85.05 74.14 88.62 97.93
[…]
RemoteCLIP 1.41 86.23 35.03 2.72 2.15 87.06 94.51 74.14 95.02 1.07 92.74 79.32 85.10 71.18 87.98 97.31
[…]
Scale-MAE 1.48 86.56 34.77 2.75 2.16 88.34 95.52 69.08 94.68 1.10 94.94 77.21 86.11 68.18 89.69 96.40
[…]
SSL4EO-S12 1.51 82.29 34.90 2.89 2.34 83.30 95.87 69.16 93.65 0.96 91.45 70.28 88.89 64.73 87.12 95.62
[…]
ViT 1.46 77.82 34.74 2.86 2.20 74.31 90.58 59.93 93.17 1.10 92.73 74.52 79.79 34.67 82.71 93.17
[…]
We freeze the FM encoder weights and fine-tune only the decoder parameters [29]. FMs, with their pretrained encoders frozen and decoders fine-tuned on SDG-specific tasks, generally exhibit superior performance across diverse tasks than trained-from-scratch traditional models. This advantage comes from their ability to leverage large-scale pretraining, allowing them to adapt and generalize more effectively across various SDG challenges. Additionally, based on our experiments, FMs converge faster in the training stage than traditional models, as their pretraining will enable them to start with established knowledge. This advantage helps to reduce the time and data needed for fine-tuning on SDG tasks. On the other hand, traditional models like ViT and ResNet-50, which are trained from scratch for each SDG task, may require more extensive tuning to achieve optimal performance, especially for broader generalization. While FMs often excel in flexibility, faster convergence, and achieving state-of-the-art results, traditional models can still perform competitively on a few tasks.
Model Finetuning Decoder Only Finetuning Full Parameters
mF1 $\uparrow$ Training Energy (kWh) Inference Energy (kWh) CO2 Emission (kg) mF1 $\uparrow$ Training Energy (kWh) Inference Energy (kWh) CO2 Emission (kg) CO2 Emission Relative Difference (%)
[…]
DOFA 89.328 0.0359 0.00029 0.02121 89.262 0.05288 0.00029 0.03111 +39.46
[…]
Prithvi 87.864 0.03473 0.00029 0.02104 87.964 0.05246 0.00028 0.03071 +61.11
[…]
SatlasNet 87.416 0.03428 0.00041 0.02028 89.19 0.09296 0.0004 0.05436 +168.05
[…]
SpectralGPT 87.663 0.19017 0.00142 0.11206 87.75 0.45666 0.0016 0.26864 +139.73
[…]
SoftCon 88.191 0.0417 0.00033 0.02498 88.089 0.06883 0.00041 0.04055 +62.33
Trained from Scratch for traditional deep models
[…]
ResNet-50 - - - - 82.756 0.03782 0.00011 0.02346 -
Discussions
Table 3 summarizes the strengths and weaknesses of FMs (decoder fine-tuned) compared to traditional models in terms of performance, efficiency, and sustainability criteria. Here, we discuss our key findings, as well as principles and considerations for geospatial FMs.
Assessment Criteria FMs (Decoder Fine-tuned) Traditional Models
[…]
Training Efficiency (kWh) ✓lower training cost ✓moderate
[…]
CO2 Emissions (kg) ✓generally low ✓moderate
[…]
Key Findings
Beyond the Hype
Aligning models with the nature of EO: Training FMs from scratch requires an enormous amount of time and computational resources, which make it largely impractical for benchmarking studies like ours. Given these constraints, we focused on fine-tuning existing pre-trained models. Accordingly, the energy use reported in our work reflects only this fine-tuning phase. We strongly encourage researchers to clearly report CO2 emissions or other environmental impact measures whenever they introduce new FMs to provide the whole picture. This transparency is essential to support responsible AI development and help the community move toward more sustainable practices in training and using large models.
The development of geospatial FMs demands a methodological shift that reflects the unique nature of EO data. As argued in [32], EO data is not just another visual modality; They include unique properties, such as spectral and temporal richness, geolocation, physical constraints, and sensor-specific noise patterns, which are not adequately addressed by conventional vision-based pretraining objectives such as masked image modeling or contrastive learning or the models themselves [33]. As a result, downstream geospatial tasks such as change detection, time-series forecasting, and biophysical retrieval may suffer from a misalignment between model pretraining and real-world task demands. Moreover, relying on standard computer vision architectures, such as UperNet or Siamese UperNet, can hide the real impact of task-specific reasoning in geospatial applications. To move the field forward, we advocate for models that capture the specific nature of geospatial data. This includes using physics-informed pretraining that brings in domain knowledge, learning methods that combine different types of data (like spatial, spectral, and temporal), and flexible decoder designs with suitable loss functions that match the needs of specific geospatial tasks.
Furthermore, the importance of scale must not be overlooked. Though originally developed for language models, scaling laws emphasize the importance of maintaining an appropriate data-to-parameter ratio, typically ranging from 5:1 to 20:1, to achieve generalizable performance [34]. As highlighted in [33], many geospatial FMs are undertrained relative to their size, limiting their generalization capacity despite architectural scale. Smaller models risk underfitting complex Earth patterns, while larger models require correspondingly large and diverse training datasets to fulfill their potential.
Last but not least, we want to emphasize that while technical benchmarking is a necessary step, it is not a sufficient one. Public perceptions significantly influence policy decisions by shaping political priorities, public support, and the feasibility of implementation. Performance metrics alone do not guarantee real-world relevance or societal impact. While benchmarks offer a standardized way to measure methodological progress, they must be grounded in broader, mission-driven goals. As aligned with studies such as [35], the impactful deployment of geospatial FMs, especially in support of SDG-17, relies on collaboration with stakeholders, domain experts, and decision-makers who can translate model outputs into actionable insights for "Becoming Good at AI for Good". In practice, benchmarking and modeling represent just two components of a complete solution. Defining relevant problems, selecting appropriate metrics, involving domain experts, and clearly articulating the model’s purpose are equally important. FMs, unlike narrow models that suffer from “catastrophic forgetting”, offer unique advantages for long-term monitoring and adaptation to temporal shifts in the Earth system. Well-pretrained FMs enable lifelong learning and can generalize to new tasks with minimal further tuning.
Methods
Achieving the SDGs requires reliable tools to monitor progress across complex, interconnected systems. Despite the global availability of satellite data, assessing the real-world impact of development processes and their alignment with sustainability goals remains a major challenge [2]. We introduce SustainFM, a benchmark dataset for monitoring 16 SDGs using EO data. Table 4 provides the overview of SustainFM, where each entry corresponds to a distinct task aligned with a specific SDG-related application. To ensure broad applicability, SustainFM includes data from a variety of satellite sensors, spatial resolutions, geographies, and AI task types, including regression (Reg.), classification (Cls.), semantic segmentation (Seg.), and change detection (CD), capturing the diversity of real-world use cases. Given the broad spectrum of AI tasks associated with different SDGs, we attach task-specific decoders to each FM encoder. For semantic segmentation, we use a UperNet decoder; for change detection, we use a Siamese UperNet; and for classification and regression, we use a fully connected layer with global pooling. Optimization uses cross-entropy loss for classification, segmentation, change detection tasks, and mean squared error (MSE) loss for regression.
SustainFM is designed as a testbed for downstream tasks in FMs to enable systematic evaluation of performance, generalization, and energy efficiency across data modalities and learning tasks. SustainFM includes multiple EO sources, including Landsat, Sentinel-1/2, VIIRS, Gaofen-2, Google Earth, and PlanetScope, with spatial resolutions ranging from 0.5 m to 30 m. Geographically, the dataset spans globally, covering over 200 countries and regions, to minimize regional bias. The tasks tackle real-world challenges such as health, education, infrastructure, industrial activity, and disaster response. Below, we introduce the 16 building blocks of SustainFM, each aligned with a specific SDG, along with the model evaluation protocols.
@@ Methods
2 Cropland Change Detection Gaofen-2 0.5-2m China CD $256\times 256256 × 256$ 2,400
@@
4 Women Educational Attainment Regression Landsat+VIIRS 30m Africa Reg. $256\times 256256 × 256$ 117,062
@@
6 Urban BOW Detection Gaofen-2 1m China Seg. $256\times 256256 × 256$ 1,645
@@
8 Detection of Settlements without Electricity Sent-1+Sent-2+Landsat+VIIRS 10m Malawi Seg. $800\times 800800 × 800$ 98
@@
10 Sanitation Index Regression Landsat+VIIRS 30m Africa Reg. $256\times 256256 × 256$ 89,271
@@
12 Mining Change Detection Google Earth 1.2m Global CD $256\times 256256 × 256$ 71,711
@@
14 Marine Pollution Detection Sent-2 10m Global Seg. $240\times 240240 × 240$ 2,803
@@
16 War-induced Damage Classification Planet Scope 3m Regional Cls. $128\times 128128 × 128$ 13,460
SustainFM Benchmark Dataset

### v2

@@ Abstract
Foundation Models (FMs) are large-scale, pre-trained artificial intelligence (AI) systems that have revolutionized natural language processing and computer vision, and are now advancing geospatial analysis and Earth Observation (EO). They promise improved generalization across tasks, scalability, and efficient adaptation with minimal labeled data. However, despite the rapid proliferation of geospatial FMs, their real-world utility and alignment with global sustainability goals remain underexplored. We introduce SustainFM, a comprehensive benchmarking framework grounded in the 17 Sustainable Development Goals with extremely diverse tasks ranging from asset wealth prediction to environmental hazard detection. This study provides a rigorous, interdisciplinary assessment of geospatial FMs and offers critical insights into their role in attaining sustainability goals. Our findings show: (1) While not universally superior, FMs often outperform traditional approaches across diverse tasks and datasets. (2) Evaluating FMs should go beyond accuracy to include transferability, generalization, and energy efficiency as key criteria for their responsible use. (3) FMs enable scalable, SDG-grounded solutions, offering broad utility for tackling complex sustainability challenges. Critically, we advocate for a paradigm shift from model-centric development to impact-driven deployment, and emphasize metrics such as energy efficiency, robustness to domain shifts, and ethical considerations.
@@
Keywords Foundation Models, Earth Observation, Sustainable Development Goals, Geospatial AI, Benchmarking.
@@
1 Introduction
The convergence of artificial intelligence (AI) and Earth observation (EO) technologies has launched an unprecedented era of capabilities for addressing global sustainability challenges [1]. The Sustainable Development Goals (SDGs), established by the United Nations General Assembly in 2015, represent a comprehensive framework for addressing global challenges through 17 interconnected goals spanning social, economic, and environmental dimensions [2, 3, 4]. These goals aim to achieve a better and more sustainable future by 2030, addressing critical issues including poverty, inequality, climate change, environmental degradation, peace, and justice [5, 6, 7]. EO techniques significantly contribute to sustainable development by providing critical data and insights that support the achievement of SDGs [6, 7, 8]. From monitoring climate change, biodiversity, and water resources to enhancing food security, urban planning, and disaster resilience [9, 10, 11, 12], EO provides critical information for tracking environmental and socio-economic trends. Through supporting evidence-based policy and providing an enormous amount of up-to-date data, EO accelerates progress toward a more sustainable, equitable, and resilient future [13, 14].
The scale and complexity of EO data necessitate automated analytical tools capable of handling diverse, multimodal spatio-temporal datasets [15]. Foundation models (FMs) have emerged as a transformative solution, offering unprecedented capabilities for geospatial applications through their ability to process vast amounts of heterogeneous data and generalize across multiple downstream tasks [16, 17]. FMs represent a paradigm shift in AI, characterized by large-scale pre-training on extensive, diverse datasets using self-supervised [18] or fully supervised learning approaches [19]. These models normally contain billions of parameters (e.g., Google’s SegGPT [20] with over 3 billion and Meta’s DINOv2 [21] with around 1 billion parameters), and leverage vast computational resources to learn detailed patterns and representations from massive datasets, which enables them to generalize across a wide range of downstream tasks. With this definition, they serve as a universal backbone to support a wide range of applications across domains such as natural language processing (NLP), computer vision, and (geo)scientific computing. The term foundation emphasizes its role in providing a pre-trained, transferable representation that can be fine-tuned or adapted for domain-specific applications. Specifically, geospatial FMs are pre-trained on vast amounts of airborne and spaceborne data and can then be adapted for downstream EO tasks, such as land cover classification, change detection, or disaster response mapping, without requiring separate models to be trained from scratch for each task [22]. Key characteristics of FMs include: (1) Scalability through efficient, distributed training; (2) Generalization via transferable representations across tasks; and (3) Few-/Zero-shot Learning, enabling adaptation with little or no labeled data.
@@ 1 Introduction
The evolution of geospatial FMs has witnessed explosive growth, progressing from single-modal approaches focusing on optical satellite imagery [23, 24, 25] to sophisticated multi-modal architectures that integrate diverse data sources [26, 27, 28, 29], including optical, SAR, hyperspectral, and auxiliary geospatial information [30, 31]. However, the rapid growth in the number of FMs over such a short period has led to a saturation of new architectures, resembling the clear trajectory we witnessed in the EO community during the deep learning boom around 2016, often marked by incremental improvements and limited consideration of their societal, environmental, and economical impact. A fundamental gap remains in grounding model development with global sustainability objectives. To see the actual impact of FMs on real-world challenges, such as climate monitoring, resource management, and humanitarian applications, there needs to be a stronger focus on the deployment phase (downstream tasks), which is widely regarded as the ultimate goal of Responsible AI in EO [32].
On the other hand, current evaluation approaches primarily focus on task-specific accuracy metrics [33, 34, 35], providing limited insight into the model’s utility for addressing complex sustainability challenges. This gap is particularly concerning when FMs are increasingly applied in domains directly related to human welfare, environmental conservation, and economic development areas, where decisions carry profound consequences beyond accuracy scores. To truly unlock their potential, FMs should be assessed not only on technical sophistication but also on their contributions to social equity, environmental protection, and economic sustainability. Yet, such dimensions are often absent in current evaluation frameworks, underscoring the urgent need for a systematic benchmark that integrates mainstream FMs with multi-source datasets aligned to the SDGs.
@@
To address this, we must shift the emphasis from mere model development to the deployment phase [36], where FMs are evaluated based on their impact on complex socio-environmental contexts. Performance evaluation should go beyond traditional benchmarks and include real-world scenarios, social good, and ethical considerations in the deployment stage. Furthermore, the interdisciplinary views of both AI and domain experts are required to ensure that FMs are not only optimized for accuracy but also for their contribution to long-term societal benefits [37]. Without this shift, we risk an unsustainable model development cycle with no real-world impacts.
@@ 1. Do we need FMs, or are traditional, purpose-specific deep learning models sufficient for EO tasks?
Through systematic evaluation across key criteria, including accuracy, transferability, data efficiency, energy costs, and operational impact, we aim to provide a grounded, rigorous assessment that challenges assumptions and clarifies the actual added value, if any, of FMs in this domain. Ultimately, our goal is to contribute to global sustainability objectives and promote social good, a key aspect that has been undermined in many current geospatial FMs. To evaluate FMs in EO rigorously, we introduce SustainFM, a benchmark aligned with all 17 SDGs that provides new EO datasets, leverages pretrained FMs, and evaluates them for global sustainability, as conveyed in Fig. 1.
This paper makes four key contributions. Scientifically, it introduces SustainFM, the first comprehensive benchmarking framework for evaluating FMs against the SDGs, setting new standards for responsible AI in Earth observation. Methodologically, it advances beyond accuracy-focused assessments by integrating critical metrics such as energy efficiency, transferability, and environmental impact, promoting a paradigm shift from performance-driven development to impact-driven deployment. Societally, it demonstrates how FMs can enable scalable solutions to global challenges while ensuring equitable outcomes across diverse communities. Environmentally, it quantifies the carbon footprint of model training and advocates for energy-efficient strategies that reduce emissions without sacrificing performance. By aligning geospatial AI with the SDG framework, this work lays a crucial foundation for leveraging advanced AI responsibly to address pressing sustainability challenges.
@@
2 Cropland Change Detection Gaofen-2 0.5-2m China CD $256\times 256$ 2,400
@@
4 Women Educational Attainment Regression Landsat+VIIRS 30m Africa Reg. $256\times 256$ 117,062
@@
6 Urban BOW Detection Gaofen-2 1m China Seg. $256\times 256$ 1,645
@@
8 Detection of Settlements without Electricity Sent-1+Sent-2+Landsat+VIIRS 10m Malawi Seg. $800\times 800$ 98
@@
10 Sanitation Index Regression Landsat+VIIRS 30m Africa Reg. $256\times 256$ 89,271
@@
12 Mining Change Detection Google Earth 1.2m Global CD $256\times 256$ 71,711
@@
14 Marine Pollution Detection Sent-2 10m Global Seg. $240\times 240$ 2,803
@@
16 War-induced Damage Classification Planet Scope 3m Regional Cls. $128\times 128$ 13,460
2 SustainFM
Achieving the SDGs requires reliable tools to monitor progress across complex, interconnected systems. Despite the global availability of satellite data, assessing the real-world impact of development processes and their alignment with sustainability goals remains a major challenge [7]. To address this gap, we introduce SustainFM, a benchmark dataset for monitoring $16$ SDGs using EO data. Specifically, SustainFM spans over 200 countries and regions across continents (Fig. 2), covering diverse sensor types and spatial resolutions ($0.5$ m–$30$ m), including Landsat, Sentinel-1/2, VIIRS, Gaofen-2, HLS, Google Earth, and PlanetScope, to reflect the complexity of global sustainability challenges. It integrates 16 real-world tasks aligned with the first 16 SDGs, including asset wealth prediction, health estimation, solar farm and settlement mapping, as well as detection of industrial smoke, marine pollution, and conflict-related damage. These tasks represent key learning paradigms, including regression (Reg.), classification (Cls.), semantic segmentation (Seg.), and change detection (CD). Through collaboration with experts from diverse fields, including remote sensing, machine learning, statistics, and AI ethics, SustainFM also contributes to SDG-17 (Partnerships for the Goals), bridging the gap between academic benchmarks and real-world impact. Table 1 summarizes the EO datasets in SustainFM, with each entry corresponding to a task linked to a specific SDG application. Designed as a testbed for downstream tasks, SustainFM supports systematic evaluation of FM performance, generalization, and energy efficiency across data modalities and learning settings. Below, we introduce the 16 building blocks of SustainFM, each aligned with a specific SDG, along with the corresponding evaluation protocols.
2.1 No Poverty (SDG-1)
Eradicating poverty in all its forms is a central objective of sustainable development. Effective monitoring of poverty requires capturing spatial and socioeconomic disparities, which can be challenging using traditional survey methods alone. EO data offers a complementary approach by providing large-scale insights into analyzing poverty triggers from a bird’s-eye view of residential conditions, including human settlements, infrastructure quality, and resource distribution [38, 39, 40]. However, human development circumstances within society, such as human well-being, education, and gender equality, still face difficulties in being directly measured or indicated by EO data due to the lack of labels.
To tackle these challenges, recent literature has seen innovations that indicate the development of such themes by using human survey data and subsequently connecting these survey data with satellite images via regression tasks [41, 42]. One possible solution is to use the Demographic and Health Surveys (DHS) survey data, which has been widely explored for its relationship with the satellite images. For example, SustainBench [43] explored six variables in the DHS dataset, including asset wealth index, women’s BMI, child mortality rate, women’s educational attainment, water quality index, and sanitation index. In particular, the SustainBench processes the DHS survey data by converging the localized entries into cluster-level data, and subsequently collects satellite images that are geographically centered around each surveyed cluster’s geocoordinates. These satellite images comprise both daytime images from the Landsat 5, 7, and 8 satellites and nightlight images from the DMSP and VIIRS satellites.